OpenStudy (anonymous):

Given the following endpoints (A & B) and midpoint (M) on a number line, find the coordinates for A & B. M(-11) A(2x + 3) B(5x + 3) 1) A(-5) & B(-17) 2) A(-9) & B(-13) 3) A(-13) & B(-9) 4) A(-17) & B(-5)

OpenStudy (amistre64):

(A+B) ----- = M 2

OpenStudy (amistre64):

(2x+3+5x+3)/2 = -11 7x +6 = -22 7x = -28 x = -4 maybe?

OpenStudy (amistre64):

my guess is A